OpenStudy (anonymous):

A keycode must contain 2 letters and 3 numbers. The letters may be any letter of the alphabet. The numbers should be any number from 0 to 9. How many different keycode combinations are there? A. 712,000 B. 676,000 C. 852,000 D. 1,576,000

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello

OpenStudy (misty1212):

use the "counting principle" i.e multiply all the choices together

OpenStudy (misty1212):

26 choices for each letter, 10 for each number, do \[26\times 26\times 10\times 10\times 10\]

OpenStudy (anonymous):

thank you <3

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]